OpenStudy (anonymous):

Yolanda invests her $5000 bonus into a high yield account that earns 6.7% simple interest. She wants to buy a Burmese Rabbit Hound with the money. The breeder told her that the puppy would cost $5500. How long will Yolanda have to wait before she has $5500 in her savings account? Is this right? t=I/Pr t=5500/(5000)(0.067) t=5500/335 t=16.42

OpenStudy (anonymous):

The formula is correct. However I = 500, not 5500. Yolanda besides her $5000, she needs $500 earning from the interest to have a total of $5500. The interest she can make in one year is Pr where P = 5000 and r = 0.067 (r must be in decimal form. For example, 3% is 0.03.) Pr = 5000 x 0.067 = 335 since she needs $500, we can see that she will have to wait a little longer than a year. Since I = Prt => t = I/Pr = 500/335 = 1.49 ~ 1.5 years or 18 months. When solving real life problems, please use your common sense. Yolanda needs 1/10 of her money. Since she can earn 6.7% or $6.70 for every $100 she has, she can earn approximately more than $300 per year. She doesn't have to wait more than 16 years! Many students while solving math try to memorize the formulas then use the given numbers to plug them in without even thinking what is going on. The purpose of learning math is to solve your daily life problem logically, not to just get the correct answers so you can have better grade. When you understand the problem and know how to solve them, the best grade will follow. :)

OpenStudy (anonymous):

Thank you, I'll try to remember all that while solving problems next time:)